Tohban Report 2014-10-08
From RHESSI Wiki
|Start Date:||01 Oct 2014|
|End Date:||08 Oct 2014|
|Tohban:||Juan Camilo Buitrago-Casas|
|Next Tohban:||Juan Carlos Martínez-Oliveros|
|List all reports|
Activity has been significantly low this week though a very small active region (AR-2178) was making its rounds on the disk. It will soon go around the west limb. Two M-class flares occurred on October 2. A medium size active region is peeking through the East.
SSR is well-behaved though a little high (~35%) during October 1 and October 2. The decimation was changed to "Active/Vigorous" at 276-00:46 UTC. Additionally, David's lightning experiment is still ongoing well. Some extra passes have been added from Weilheim and Santiago.
Two warnings have appeared regarding the Cold plate 1 [125.52 degrees (K)] and the Solar Array Temperature 1 [97.698 degrees (C)].
Normal as far as a tohban can tell. Rhessi was spun up to 15 RPM on October 4 between 00:30 and 02:38 UTC (October 2, 17:30-19:38 Berkeley time). Attenuator motion was nominal.
No significant data gaps.
Detector 2: We still have the light leak problem though it is sort of stable. Other detectors have seemed ok with light leak so far.
Detector 9: It showed some periodical spikes during last couple of days. Next Tohban should kept an eye on that issue.
There were no threshold changes this week.
The SDO roll manoeuvre is being conducted from 2014/281 (October 8) 17:51 to 282 (October 9) 00:41 during RHESSI is collecting as much RHESSI/SAS data as possible. With this in mind, the ATS has been switched. The special ADP commanding will run at 2014/281-17:50 UTC, and will conclude at 2014/282-00:47 UTC.
The hsi_tohban_flare_counter.pro code showed a problem with data from GOES.
No GOES flares for Time Range
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||full nighttime events|
|Night time data (rears)||full nighttime events|
|Detector problems?||Detector 9 spikes|